OpenStudy (anonymous):

Rite cut riding lawnmower obeys the demand equation p=-1/20x+1060. The cost of producing x lawnmowers is given by the function c(x)=120x+5000 a)Express the revenue R as a function of x b)Express the profit P as a function of X c)Fine the value of x that maximizes profit. What is the maximum profit? d)What price should be charged to maximize profit?

OpenStudy (anonymous):

R = P(x)*x = (-1/20)+1060x P = R - C = (-1/20)+1060x - (120x+5000)

OpenStudy (anonymous):

Assuming perfect competition p = MC = 120 Plug that into your inverse demand to get quantity